25 math wizards are participating in math contest . in how many ways can there be a first, second and third places?
Answer by stanbon(75887)

25 math wizards are participating in math contest . in how many ways can there be a first, second and third places?
----
Order makes a difference, so 25P3 = 25!/(25-3)! = 25*24*23 = 13,800 ways
---------------
Cheers,
Stan H.
----------